# Distributional Bellman Operators over Mean Embeddings

### 09 Dec 2023 v1

@@ Abstract
We propose a novel algorithmic framework for distributional reinforcement learning, based on learning finite-dimensional mean embeddings of return distributions. We derive several new algorithms for dynamic programming and temporal-difference learning based on this framework, provide asymptotic convergence theory, and examine the empirical performance of the algorithms on a suite of tabular tasks. Further, we show that this approach can be straightforwardly combined with deep reinforcement learning, and obtain a new deep RL agent that improves over baseline distributional approaches on the Arcade Learning Environment.
@@ 1 Introduction
In distributional approaches to reinforcement learning (RL), the aim is to learn the full probability distribution of future returns (Morimura et al., 2010a; Bellemare et al., 2017; 2023), rather than just their expected value, as is typically the case in value-based reinforcement learning (Sutton & Barto, 2018). Distributional RL was proposed in the setting of deep reinforcement learning by Bellemare et al. (2017), with a variety of precursor work stretching back almost as far as Markov decision processes themselves (Jaquette, 1973; Sobel, 1982; Chung & Sobel, 1987; Morimura et al., 2010a; b). Beginning with the work in Bellemare et al. (2017), the distributional approach to reinforcement learning has been central across a variety of applications of deep RL in simulation and in the real world (Bodnar et al., 2020; Bellemare et al., 2020; Wurman et al., 2022; Fawzi et al., 2022).
@@
Typically, predictions of return distributions are represented directly as approximate probability distributions, such as categorical distributions (Bellemare et al., 2017). Rowland et al. (2019) proposed an alternative framework where return distributions are represented via the values of statistical functionals, called a sketch by Bellemare et al. (2023). This provided a new space of distributional reinforcement learning algorithms, leading to improvements in deep RL agents, and hypotheses regarding distributional RL in the brain (Dabney et al., 2020; Lowet et al., 2020). On the other hand, a potential drawback of this approach is that each distributional Bellman update to the representation, these values must be “decoded” back into an approximate distribution via an imputation strategy. In practice, this can introduce significant computational overhead to Bellman updates, and is unlikely to be biologically plausible for distributional learning in the brain (Tano et al., 2020).
@@
Here, we focus on a notable instance of the sketch called the mean embedding sketch. In short, the mean embedding is the expectation of nonlinear functions under the distribution represented (Smola et al., 2007; Sriperumbudur et al., 2010; Berlinet & Thomas-Agnan, 2011), and is related to frames in signal processing (Mallat, 1999) and distributed distributional code in neuroscience (Sahani & Dayan, 2003; Vértes & Sahani, 2018). The core contributions of this paper are to revisit the approach to distributional reinforcement learning based on sketches (Rowland et al., 2019), and to propose the sketch Bellman operator that updates the implicit distributional representation as a simple linear operation, obviating the need for the expensive imputation strategies converting between sketches and distributions. This provides a rich new space of distributional RL algorithms that operate entirely in the space of sketches. We provide theoretical convergence analysis to accompany the framework, investigate the practical behaviour of various instantiations of the proposed algorithms in tabular domains, and demonstrate the effectiveness of the sketch framework in deep reinforcement learning, showing that our approach is robust enough to serve as the basis for a new variety of deep distributional reinforcement learning algorithms.
@@ 2 Background
We consider a Markov decision process (MDP) with state space $\mathcal{X}$, action space $\mathcal{A}$, transition probabilities $P:\mathcal{X}\times\mathcal{A}\rightarrow\mathscr{P}(\mathcal{X})$, reward distribution function $P_{R}:\mathcal{X}\times\mathcal{A}\rightarrow\mathscr{P}(\mathbb{R})$, and discount factor $\gamma\in[0,1)$. Given a policy $\pi:\mathcal{X}\rightarrow\mathscr{P}(\mathcal{A})$ and initial state $x\in\mathcal{X}$, a random trajectory $(X_{t},A_{t},R_{t})_{t\geq 0}$ is the sequence of random states, actions, and rewards encountered when using the policy $\pi$ to select actions in this MDP. More precisely, we have $X_{0}=x$, $A_{t}\sim\pi(\cdot|X_{t})$, $R_{t}\sim P_{R}(X_{t},A_{t})$, $X_{t+1}\sim P(\cdot|X_{t},A_{t})$ for all $t\geq 0$. We write $\mathbb{P}^{\pi}_{x}$ and $\mathbb{E}^{\pi}_{x}$ for probabilities and expectations with respect to this distribution, respectively. The performance along the trajectory is measured by the discounted return, defined by
@@
In typical value-based reinforcement learning, during policy evaluation, the agent learns the expectation of the return for each possible initial state $x\in\mathcal{X}$, which is encoded by the value function $V^{\pi}:\mathcal{X}\rightarrow\mathbb{R}$, given by $V^{\pi}(x)=\mathbb{E}^{\pi}_{x}[\sum_{t=0}^{\infty}\gamma^{t}R_{t}]$.
2.1 Distributional RL and the distributional Bellman equation
In distributional reinforcement learning, the problem of policy evaluation is to learn the probability distribution of return in Equation (1) for each possible initial state $x\in\mathcal{X}$. This is encoded by the return-distribution function $\eta^{\pi}:\mathcal{X}\rightarrow\mathscr{P}(\mathbb{R})$, which maps each initial state $x\in\mathcal{X}$ to the corresponding distribution of the random return, e.g. $\eta^{\pi}(x)$ is the return distribution of state $x$. A central result in distributional reinforcement learning is the distributional Bellman equation, which relates the distribution of the random return under different combinations of initial states and actions.
To build the random variable formulation of the returns, we let $(G^{\pi}(x):x\in\mathcal{X})$ be a collection of random variables with the property that $G^{\pi}(x)$ is equal to Equation (1) in distribution, conditioned on the initial state $X_{0}=x$. This formulation implies that the random variable $G^{\pi}(x)$ is distributed as $\eta^{\pi}(x)$, introduced above, for all $x\in\mathcal{X}$. Consider a random transition $(x,R,X^{\prime})$ generated by $\pi$, independent of the $G^{\pi}$ random variables. Then, the (random variable) distributional Bellman equation states that for each initial state $x$,
@@ 2.1 Distributional RL and the distributional Bellman equation
Here, we use the slight abuse of the conditioning bar to set the distribution of $X$ in the random transition. It is also useful to introduce the distributional Bellman operator $\mathcal{T}^{\pi}:\mathscr{P}(\mathbb{R})^{\mathcal{X}}\rightarrow\mathscr{P}(\mathbb{R})^{\mathcal{X}}$ to describe the transformation that occurs on the right-hans side (Morimura et al., 2010a; Bellemare et al., 2017). If $\eta\in\mathscr{P}(\mathbb{R})^{\mathcal{X}}$ is a collection of probability distributions, and $(G(x):x\in\mathcal{X})$ is a collection of random variables such that $G(x)\sim\eta(x)$ for all $x$, and $(X,R,X^{\prime})$ is random transition generated by $\pi$, independent of $(G(x):x\in\mathcal{X})$, then $(\mathcal{T}^{\pi}\eta)(x)=\text{Dist}(R+\gamma G(X^{\prime})|X=x)$.
@@ 2.2 Statistical functionals and sketches
For a given distribution $\nu$, the embedding $\psi(\nu)$ can therefore be thought of as providing a lossy summary of the distribution. The name here is motivated by the kernel literature, in which Equation (2) can be viewed as embedding the distribution $\nu$ into $\mathbb{R}^{m}$ based on the mean value of $\phi$ under this distribution (Smola et al., 2007; Sriperumbudur et al., 2010; Berlinet & Thomas-Agnan, 2011).
Figure 1: The statistical functional framework proposed by Rowland et al. (2019) (top), and the framework proposed in this paper, directly updating sketches, avoiding the imputation step (bottom).
Statistical functional dynamic programming and temporal-difference learning (SFDP/SFTD; Rowland et al. (2019), see also Bellemare et al. (2023)) is an approach to distributional RL in which sketch values, rather than approximate distributions, are the primary object learned. Given a sketch $\psi$ and estimated sketch values $U:\mathcal{X}\rightarrow\mathbb{R}^{m}$, these approaches proceed by first defining an imputation strategy $\iota:\mathbb{R}^{m}\rightarrow\mathscr{P}(\mathbb{R})$ mapping sketch values back to distributions, with the aim that $\psi(\iota(U))\approx U$, so that $\iota$ behaves as an approximate pseudo-inverse to $\psi$. The usual Bellman backup is then applied to this imputed distribution, and the sketch value extracted from this updated distribution. Thus, if $U:\mathcal{X}\rightarrow\mathbb{R}^{m}$ represents approximations to sketch values, a typical update in SFDP takes the form $U\leftarrow\psi((\mathcal{T}^{\pi}\iota(U))(x))$ (Bellemare et al., 2023); see Figure 1.
This approach led to expectile-regression DQN, a deep RL agent that aims to learn the sketch values associated with certain expectiles (Newey & Powell, 1987) of the return, and influenced a distributional model of dopamine signalling in the brain (Dabney et al., 2020). An important consideration is that computation of the imputation strategy is often costly in machine learning applications, and considered biologically implausible in neuroscience models (Tano et al., 2020).
@@ 3 The Bellman sketch framework
Our goal is to derive a framework for approximate computation of the sketch $\psi$ (with corresponding feature function $\phi$) of the return distributions corresponding to a policy $\pi$, without needing to design, implement, and compute an imputation strategy as in the case of SFDP/TD; see Figure 1 for a visual comparison of the two approaches. That is, we aim to compute the object $U^{\pi}:\mathcal{X}\rightarrow\mathbb{R}^{m}$, given by
@@
We begin by considering environments with a finite set of possible rewards $\mathcal{R}\subseteq\mathbb{R}$; we discuss generalisations later. To motivate our method, we first consider a special case; suppose that for each possible return $g\in\mathbb{R}$, and each possible immediate reward $r\!\in\!\mathcal{R}$, there exists a matrix $B_{r}$ such that
@@
$U^{\pi}(x)~{}\overset{(a)}{=}~{}\mathbb{E}^{\pi}_{x}[\phi(R+\gamma G^{\pi}(X^{\prime}))]~{}\overset{(b)}{=}~{}\mathbb{E}^{\pi}_{x}[B_{R}\phi(G^{\pi}(X^{\prime}))]~{}\overset{(c)}{=}~{}\mathbb{E}^{\pi}_{x}[B_{R}U^{\pi}(X^{\prime})]\,,$ (4)
@@
As this is an update applied directly to sketch values themselves, we introduce the sketch Bellman operator $\mathcal{T}^{\pi}_{\phi}:(\mathbb{R}^{m})^{\mathcal{X}}\rightarrow(\mathbb{R}^{m})^{\mathcal{X}}$, with $(\mathcal{T}^{\pi}_{\phi}U)(x)$ defined according to the right-hand side of Equation (Sketch-DP). Note that $\mathcal{T}^{\pi}_{\phi}$ is a linear operator, in contrast to the standard expected-value Bellman operator, which is affine. We recover the affine case by taking one component of $\phi$ to be constant, e.g. $\phi_{1}(g)\equiv 1$, and enforcing $U_{1}(x)\equiv 1$.
The right-hand side of Equation (Sketch-DP) can be unbiasedly approximated with a sample transition $(x,r,x^{\prime})$. Stochastic approximation theory (Kushner & Yin, 1997; Bertsekas & Tsitsiklis, 1996) then naturally suggests the following temporal-difference learning update, given a learning rate $\alpha$:
$U(x)~{}\leftarrow~{}(1-\alpha)U(x)+\alpha B_{r}V(x^{\prime})\,,$ (Sketch-TD)
Rowland et al. (2019) introduced the term Bellman closed for sketches for which an exact dynamic programming algorithm is available, and provided a characterisation of Bellman closed mean embedding sketches. The notion of Bellman closedness is closely related to the relationship in Equation (3), and from Rowland et al. (Theorem 4.3; 2019), we can deduce that the only mean embedding sketches that satisfy Equation (3) are invertible linear combinations of first-$m$ moments.
@@ 3.1 General sketches
$U^{\pi}(x)~{}\overset{(a)}{=}~{}\mathbb{E}^{\pi}_{x}[\phi(R+\gamma G^{\pi}(X^{\prime}))]~{}\approx~{}\mathbb{E}^{\pi}_{x}[B_{R}\phi(G^{\pi}(X^{\prime}))]~{}\overset{(c)}{=}~{}\mathbb{E}^{\pi}_{x}[B_{R}U^{\pi}(X^{\prime})]\,,$ (6)
@@
# Precompute Bellman coefficients
Compute $C$ as in Equation (7)
for $r\in\mathcal{R}$ do
  Compute $C_{r}$ as in Equation (7)
  Set $B_{r}=C_{r}C^{-1}$
end for
Initialise $U:\mathcal{X}\rightarrow\mathbb{R}^{m}$
# Main loop
for $k=1,2,\ldots$ do
  if DP then
   $U(x)\leftarrow\!\sum\limits_{r,x^{\prime},a}P(r,x^{\prime}|x,a)\pi(a|x)B_{r}U(x^{\prime})\ \ \forall x$
  else if TD then
   Observe transition $(x_{k},a_{k},r_{k},x^{\prime}_{k})$.
   $U(x_{k})\leftarrow(1-\alpha_{k})U(x_{k})+\alpha_{k}B_{r_{k}}U(x^{\prime}_{k})$
  end if
end for
Computing Bellman coefficients. Under mild conditions (invertibility of $C$ as follows) the matrix of Bellman coefficients $B_{r}$ defined in Equation (5) can be expressed as $B_{r}=C_{r}C^{-1}$, where $C,C_{r}\in\mathbb{R}^{m\times m}$ are defined by
@@
The elements of these matrices are expressible as integrals over the real line, and hence several possibilities are available for (approximate) computation: if $\mu$ is finitely-supported, direct summation is possible; in certain cases the integrals may be analytically available, and otherwise numerical integration can be performed. Additionally, for certain feature maps $\phi$, the Bellman coefficients $B_{r}$ have particular structure that can be exploited computationally; see Appendix B.3 for further discussion. Detailed properties of $B_{r}$ are studied in Section B.5.
@@
Selecting feature maps. A natural question is what effects the choice of feature map $\phi$ has on the performance of the algorithm. There are several competing concerns. First, the richer the map $\phi$, the more information about the return distribution can be captured by the corresponding mean embedding. However, the computational costs (both in time and memory) of our proposed algorithms scale in the worst case cubically with $m$, the dimensionality of the mean embedding. In addition, the accuracy of the algorithm in approximating the mean embeddings of the true return distributions relies on having a low approximation error in Equation (6), which in turn relies on a low regression error in Equation (5) (see Proposition 4.1 below). Selecting an appropriate feature map is therefore somewhat nuanced, and involves trading off a variety of computational and approximation concerns.
A collection of feature maps we will use throughout the paper that offer the potential for trade-offs along the dimensions identified above is given by the translation family
@@
where $\kappa:\mathbb{R}\to\mathbb{R}$ is a base feature function, $s\in\mathbb{R}^{+}$ is the slope, and the set $\{z_{1},\ldots,z_{m}\}\subseteq\mathbb{R}$ is the anchors of the feature map. We will often take $\kappa$ to be commonly used bounded and smooth nonlinear functions, such as the Gaussian or the sigmoid functions, and spread the anchor points over the return range. We emphasise that in principle there are no restrictions on the feature maps that can be considered in the framework; see Section B.2 for other possible choices.
@@
Given the $m$-dimensional function space obtained from the span of the coordinate functions $\phi_{1},\ldots,\phi_{m}$, the algorithms proposed above are essentially independent of the choice of basis for this space. For any invertible matrix $M\in\mathbb{R}^{m\times m}$, replacing $\phi$ by $M^{-1}\phi$, and also $||\cdot||_{2}$ by $||\cdot||_{M^{\top}M}$ in Equation (5) gives an equivalent algorithm.
@@
It is tempting to try and obtain a more general framework by allowing non-linear regression of $\phi(r+\gamma g)$ on $\phi(g)$ in Equation (5), to obtain a more accurate fit, for example fitting a function $H:\mathbb{R}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ so that $\phi(r+\gamma g)\approx H(r,\phi(g))$. The issue is that if $H$ is not linear in the second argument, then generally $\mathbb{E}[H(r,\phi(G(X^{\prime})))]\not=H(r,\mathbb{E}[\phi(G(X^{\prime}))])$, and so step (c) in Equation (6) is not valid. However, there may be settings where it is desirable to learn such a function $H$, to avoid online computation of Bellman coefficients every time a new reward is encountered in TD learning.

### 04 Mar 2024 v3

@@ Abstract
We propose a novel algorithmic framework for distributional reinforcement learning, based on learning finite-dimensional mean embeddings of return distributions. The framework reveals a wide variety of new algorithms for dynamic programming and temporal-difference algorithms that rely on the sketch Bellman operator, which updates mean embeddings with simple linear-algebraic computations. We provide asymptotic convergence theory, and examine the empirical performance of the algorithms on a suite of tabular tasks. Further, we show that this approach can be straightforwardly combined with deep reinforcement learning.
Machine Learning, ICML
[…]
@@ 1 Introduction
In distributional approaches to reinforcement learning (RL), the aim is to learn the full probability distribution of future returns (Morimura et al., 2010a; Bellemare et al., 2017, 2023), rather than just their expected value, as is typically the case in value-based reinforcement learning (Sutton & Barto, 2018). Distributional RL was proposed in the setting of deep reinforcement learning by Bellemare et al. (2017), with a variety of precursor work stretching back almost as far as Markov decision processes themselves (Jaquette, 1973; Sobel, 1982; Chung & Sobel, 1987; Morimura et al., 2010a, b). Beginning with the work in Bellemare et al. (2017), the distributional approach to reinforcement learning has been central across a variety of applications of deep RL in simulation and in the real world (Bodnar et al., 2020; Bellemare et al., 2020; Wurman et al., 2022; Fawzi et al., 2022).
@@
Typically, predictions of return distributions are represented directly as approximate probability distributions, such as categorical distributions (Bellemare et al., 2017). Rowland et al. (2019) proposed an alternative framework where return distributions are represented via the values of statistical functionals, called a sketch by (Bellemare et al., 2023). This provided a new space of distributional reinforcement learning algorithms, leading to improvements in deep RL agents, and hypotheses regarding distributional RL in the brain (Dabney et al., 2020; Lowet et al., 2020). On the other hand, a potential drawback of this approach is that each distributional Bellman update to the representation, these values must be “decoded” back into an approximate distribution via an imputation strategy. In practice, this can introduce significant computational overhead to Bellman updates, and is unlikely to be biologically plausible for distributional learning in the brain (Tano et al., 2020).
@@
Here, we focus on a notable instance of the sketch called the mean embedding sketch. In short, the mean embedding is the expectation of nonlinear functions under the distribution represented (Smola et al., 2007; Sriperumbudur et al., 2010; Berlinet & Thomas-Agnan, 2011), and is related to frames in signal processing (Mallat, 1999) and distributed distributional code in theoretical neuroscience (Sahani & Dayan, 2003; Vértes & Sahani, 2018). The core contributions of this paper are to revisit the approach to distributional reinforcement learning based on sketches (Rowland et al., 2019), and to propose the sketch Bellman operator that updates the implicit distributional representation as a simple linear operation, obviating the need for the expensive imputation strategies converting between sketches and distributions. This leads to a rich new space of distributional RL algorithms that operate entirely in the space of sketches. We provide theoretical convergence analysis to accompany the framework, investigate the practical behaviour of various instantiations of the proposed algorithms in tabular domains, and demonstrate the effectiveness of the sketch framework in deep reinforcement learning.
@@ 2 Background
We consider a Markov decision process (MDP) with state space $\mathcal{X}$, action space $\mathcal{A}$, transition probabilities $P:\mathcal{X}\times\mathcal{A}\rightarrow\mathscr{P}(\mathcal{X})$, reward distribution function $P_{R}:\mathcal{X}\times\mathcal{A}\rightarrow\mathscr{P}(\mathbb{R})$, and discount factor $\gamma\in[0,1)$. Given a policy $\pi:\mathcal{X}\rightarrow\mathscr{P}(\mathcal{A})$ and initial state $x\in\mathcal{X}$, a random trajectory $(X_{t},A_{t},R_{t})_{t\geq 0}$ is the sequence of random states, actions, and rewards encountered when using the policy $\pi$ to select actions in this MDP. More precisely, we have $X_{0}=x$, $A_{t}\sim\pi(\cdot|X_{t})$, $R_{t}\sim P_{R}(X_{t},A_{t})$, $X_{t+1}\sim P(\cdot|X_{t},A_{t})$ for all $t\geq 0$. We write $\mathbb{P}^{\pi}_{x}$ and $\mathbb{E}^{\pi}_{x}$ for probabilities and expectations with respect to this distribution, respectively (conditioned on $X_{0}=x$). The performance along the trajectory is measured by the discounted return, defined by
@@
In typical value-based RL, during policy evaluation, the agent learns the expectation of the return for each possible initial state $x\in\mathcal{X}$, which is encoded by the value function $V^{\pi}:\mathcal{X}\rightarrow\mathbb{R}$, given by $V^{\pi}(x)=\mathbb{E}^{\pi}_{x}[\sum_{t=0}^{\infty}\gamma^{t}R_{t}]$.
2.1 Distributional RL and Bellman equation
In distributional RL, the problem of policy evaluation is to learn the probability distribution of return in Equation (1) for each possible initial state $x\in\mathcal{X}$. This is encoded by the return-distribution function $\eta^{\pi}:\mathcal{X}\rightarrow\mathscr{P}(\mathbb{R})$, which maps each initial state $x\in\mathcal{X}$ to the corresponding distribution of the random return. A central result in distributional reinforcement learning is the distributional Bellman equation, which relates the distribution of the random return under different combinations of initial states and actions.
To build the random variable formulation of the returns, we let $(G^{\pi}(x):x\in\mathcal{X})$ be a collection of random variables with the property that $G^{\pi}(x)$ is equal to Equation (1) in distribution, conditioned on the initial state $X_{0}=x$. This formulation implies that the random variable $G^{\pi}(x)$ is distributed as $\eta^{\pi}(x)$ for all $x\in\mathcal{X}$. Consider a random transition $(x,R,X^{\prime})$ generated by $\pi$, independent of the $G^{\pi}$ random variables. Then, the (random variable) distributional Bellman equation states that for each state $x$,
@@ 2.1 Distributional RL and Bellman equation
Here, we use the slight abuse of the conditioning bar to set the distribution of $X$ in the random transition. It is also useful to introduce the distributional Bellman operator $\mathcal{T}^{\pi}:\mathscr{P}(\mathbb{R})^{\mathcal{X}}\rightarrow\mathscr{P}(\mathbb{R})^{\mathcal{X}}$ to describe the transformation that occurs on the right-hand side (Morimura et al., 2010a; Bellemare et al., 2017). If $\eta\in\mathscr{P}(\mathbb{R})^{\mathcal{X}}$ is a collection of probability distributions, and $(G(x):x\in\mathcal{X})$ is a collection of random variables such that $G(x)\sim\eta(x)$ for all $x$, and $(X,R,X^{\prime})$ is random transition generated by $\pi$, independent of $(G(x):x\in\mathcal{X})$, then $(\mathcal{T}^{\pi}\eta)(x)=\text{Dist}(R+\gamma G(X^{\prime})|X=x)$.
@@ 2.2 Statistical functionals and sketches
For a given distribution $\nu$, the embedding $\psi(\nu)$ can therefore be thought of as providing a lossy summary of the distribution. The name is motivated by the kernel literature, in which Equation (2) can be viewed as embedding the distribution $\nu$ into $\mathbb{R}^{m}$ based on the mean of $\phi$ under $\nu$ (Smola et al., 2007; Sriperumbudur et al., 2010; Berlinet & Thomas-Agnan, 2011). As we will show, the mean embedding sketch enables elegant distributional RL algorithms.
Figure 1: Example of DP update for state 1 with child states 2 and 3 and return distributions $\eta_{2}$ and $\eta_{3}$. In the exact distributional DP, $\eta$’s are scaled and shifted by $f_{r}(g)=r+\gamma g$, and then weighted by the transition probabilities. In SFDP (Rowland et al., 2019), the map $\iota$ imputes, from initial sketch values $U$, approximate (e.g. categorical) distributions on which the distribution DP is applied, followed by evaluating the sketch $\psi$. In our approach Sketch-DP, the updates are computed in the mean embedding space, facilitated by the Bellman coefficients $B_{r}$, avoiding the imputation step.
Statistical functional dynamic programming (SFDP; Rowland et al. (2019), see also Bellemare et al. (2023)) is an approach to distributional RL in which sketch values, rather than approximate distributions, are the primary object learned. Given a sketch $\psi$ and estimated sketch values $U:\mathcal{X}\rightarrow\mathbb{R}^{m}$, SFDP proceed by first defining an imputation strategy $\iota:\mathbb{R}^{m}\rightarrow\mathscr{P}(\mathbb{R})$ mapping sketch values back to distributions, with the aim that $\psi(\iota(U))\approx U$, so that $\iota$ acts as an approximate pseudo-inverse of $\psi$. The usual Bellman backup is then applied to this imputed distribution, and the sketch value extracted from this updated distribution. Thus, a typical update in SFDP takes the form $U\leftarrow\psi((\mathcal{T}^{\pi}\iota(U))(x))$; see Figure 1.
This approach led to expectile-regression DQN, a deep RL agent that aims to learn the sketch values associated with certain expectiles (Newey & Powell, 1987) of the return, and influenced a distributional model of dopamine signalling in the brain (Dabney et al., 2020). An important consideration is that computation of the imputation strategy is often costly in applications and considered biologically implausible in neuroscience models (Tano et al., 2020).
@@ 3 The Bellman sketch framework
Our goal is to derive a framework for approximate computation of the sketch $\psi$ (with corresponding feature function $\phi$) of the return distributions corresponding to a policy $\pi$, without needing to design, implement, or compute an imputation strategy as in the case of SFDP/TD. That is, we aim to compute the object $U^{\pi}:\mathcal{X}\rightarrow\mathbb{R}^{m}$, given by
@@
We begin by considering environments with a finite set of possible rewards $\mathcal{R}\subseteq\mathbb{R}$; we discuss generalisations later. To motivate our method, we consider a special case; suppose that for each possible return $g\in\mathbb{R}$, and each possible immediate reward $r\!\in\!\mathcal{R}$, there exists a matrix $B_{r}$ such that
@@
$U^{\pi}(x)$ =$\overset{(a)}{=}\mathbb{E}^{\pi}_{x}[\phi(R+\gamma G^{\pi}(X^{\prime}))]\overset{(b)}{=}\mathbb{E}^{\pi}_{x}[B_{R}\phi(G^{\pi}(X^{\prime}))]$
=$\overset{(c)}{=}\mathbb{E}^{\pi}_{x}[B_{R}U^{\pi}(X^{\prime})]\,,$ (4)
@@
See Figure 1 for an example and comparison with SFDP. As this is an update applied directly to sketch values themselves, we introduce the sketch Bellman operator $\mathcal{T}^{\pi}_{\phi}:(\mathbb{R}^{m})^{\mathcal{X}}\rightarrow(\mathbb{R}^{m})^{\mathcal{X}}$, with $(\mathcal{T}^{\pi}_{\phi}U)(x)$ defined according to the right-hand side of Equation (Sketch-DP). Note that $\mathcal{T}^{\pi}_{\phi}$ is a linear operator, in contrast to the standard expected-value Bellman operator, which is affine. We recover the affine case by taking one component of $\phi$ to be constant, e.g. $\phi_{1}(g)\equiv 1$, and enforcing $U_{1}(x)\equiv 1$.
The right-hand side of Equation (Sketch-DP) can be unbiasedly approximated with a sample transition $(x,r,x^{\prime})$. Stochastic approximation theory (Kushner & Yin, 1997; Bertsekas & Tsitsiklis, 1996) then naturally suggests the following temporal-difference learning update:
$U(x)~{}\leftarrow~{}(1-\alpha)U(x)+\alpha B_{r}U(x^{\prime})\,$ (Sketch-TD)
given a learning rate $\alpha$. Rowland et al. (2019) introduced the term Bellman closed for sketches for which an exact dynamic programming algorithm is available, and provided a characterisation of Bellman closed mean embedding sketches. The notion of Bellman closedness is closely related to the relationship in Equation (3), and from Rowland et al. (Theorem 4.3; 2019), we can deduce that the only mean embedding sketches that satisfy Equation (3) are invertible linear combinations of the first-$m$ moments.
@@ 3.1 General sketches
$U^{\pi}(x)$ =$\overset{(a)}{=}\mathbb{E}^{\pi}_{x}[\phi(R+\gamma G^{\pi}(X^{\prime}))]\approx\mathbb{E}^{\pi}_{x}[B_{R}\phi(G^{\pi}(X^{\prime}))]$
=$\overset{(c)}{=}\mathbb{E}^{\pi}_{x}[B_{R}U^{\pi}(X^{\prime})]\,,$ (6)
@@
[…]
Computing Bellman coefficients. Under mild conditions (invertibility of $C$ as follows) the matrix of Bellman coefficients $B_{r}$ defined in Equation (5) can be solved as $B_{r}=C_{r}C^{-1}$, where $C,C_{r}\in\mathbb{R}^{m\times m}$ are defined by
@@
A derivation is in Section B.2 where we also describe the choice of $\mu$. The elements of these matrices are expressible as integrals over the real line, and hence several possibilities are available for (approximate) computation: if $\mu$ is finitely-supported, direct summation is possible; in certain cases the integrals may be analytically available, and otherwise numerical integration can be performed. Additionally, for certain feature maps $\phi$, the Bellman coefficients $B_{r}$ have particular structure that can be exploited computationally; see Appendix B.3 for further discussion. The generalisation to handle an infinite $\mathcal{R}$ is presented in Section B.4; and detailed properties of $B_{r}$ are studied in Section B.5.
@@
Selecting feature maps. A natural question is what effect the choice of feature map $\phi$ has on the performance of the algorithm. There are several competing concerns. First, the richer the map $\phi$, the more information about the return distribution can be captured by the corresponding mean embedding. However, in the worst case, the computational costs of our proposed Algorithm 1 scale cubically with $m$ (the dimensionality of $\phi$) prior to the iterative updates which then scale quadratically with $m$. In addition, the accuracy of the algorithm in approximating the mean embeddings of the true return distributions relies on having a low approximation error in Equation (6), which in turn relies on a low regression error in Equation (5) (see Proposition 4.1 below). Selecting an appropriate feature map is therefore somewhat nuanced, and involves trading off a variety of computational and approximation concerns.
A collection of feature maps that offer the potential for trade-offs along the dimensions above is the translation family
@@
where $\kappa:\mathbb{R}\to\mathbb{R}$ is a base feature function, $s\in\mathbb{R}^{+}$ is the slope, and the set $\{z_{1},\ldots,z_{m}\}\subseteq\mathbb{R}$ is the anchors of the feature map. We will often take $\kappa$ to be commonly used bounded and smooth nonlinear functions, such as the Gaussian or the sigmoid functions, and spread the anchor points over the return range. We emphasise that the choice of feature maps for the Bellman sketch framework is flexible; see Section B.6 for other possible choices.
@@
Given the $m$-dimensional function space obtained from the span of the coordinate functions $\phi_{1},\ldots,\phi_{m}$, the algorithms above are essentially independent of the choice of basis for this space. For any invertible matrix $M\in\mathbb{R}^{m\times m}$, replacing $\phi$ by $M^{-1}\phi$, and also $||\cdot||_{2}$ by $||\cdot||_{M^{\top}M}$ in Equation (5) gives an equivalent algorithm.
@@
It is tempting to try and obtain a more general framework by allowing non-linear regression of $\phi(r+\gamma g)$ on $\phi(g)$ in Equation (5), to obtain a more accurate fit, for example fitting a function $H:\mathbb{R}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{m}$ so that $\phi(r+\gamma g)\approx H(r,\phi(g))$. The issue is that if $H$ is not linear in the second argument, then generally $\mathbb{E}[H(r,\phi(G(X^{\prime})))]\not=H(r,\mathbb{E}[\phi(G(X^{\prime}))])$, and so step (c) in Equation (6) is not valid. However, there may be settings where it is desirable to learn such a function $H$, to avoid online computation of Bellman coefficients every time a new reward value is encountered in TD learning.
Remark 3.3 (Linear update).
The sketch updates in Equations Sketch-DP and Sketch-TD are linearly, which is distinct from typical particle-based distributional RL algorithms (e.g. (Dabney et al., 2018b, a; Nguyen-Tang et al., 2021), where the updates involve non-linear operations. In particular, Nguyen-Tang et al. (2021) proposed a TD algorithm for updating particle locations by decreasing a sample MMD objective (Gretton et al., 2012). However, this does not yield a dynamic programming algorithm, and Nguyen-Tang et al. (2021) do not analyse the TD algorithm; further comparisons are more clearly described in Appendix B.7.
ARefer to caption BRefer to caption
CRefer to caption
D                              ERefer to captionRefer to captionRefer to caption
Figure 2: An example run of Sketch-DP. A, The MRP considered here. B, The first 5 of $m=13$ sinusoidal feature functions $\phi$. The regression Equation (5) is performed under a densely spaced grid over the white region $[-4,4]$. C, Evolution of the estimated mean embeddings from initialisation (grey dot) onto the first two principal components. Crosses represent the ground-truth mean embeddings. D, Ground-truth return distributions (estimated by Monte-Carlo) and their categorical projections onto a regular grid. E, Imputed distributions from the mean embeddings onto the same grid for selected iterations (curves), compared against the categorical projections (stems).